### 20 Jan 2024 v1

# A short note on exponential sums, exceptional characters and primes in short intervals

Abstract.
This note is an overview of our new results on primes in short intervals $[x-x^{\alpha},x]$ under the assumption of the existence of exceptional Dirichlet characters. We get a small improvement of a 2004 result of Friedlander and Iwaniec. By using Nowak’s or our new estimation of exponential sums, we extend the previous admissible range $0.4937\leqslant\alpha\leqslant 1$ to $0.4925\leqslant\alpha\leqslant 1$.
Key words and phrases:
primes, short intervals, exponential sums, exceptional characters
2020 Mathematics Subject Classification:
11M06, 11N05, 11N37
1. Introduction
The famous Riemann Hypothesis is equivalent to the asymptotic formula
$\psi(x):=\sum_{n\leqslant x}\Lambda(n)=x+O\left(x^{\frac{1}{2}+\varepsilon}\right),$ (1)
where $\Lambda(n)$ is the von Mangoldt function. This implies that
$\psi(x)-\psi(x-x^{\alpha})=x^{\alpha}+O\left(x^{\frac{1}{2}+\varepsilon}\right)$ (2)
for $\frac{1}{2}+\varepsilon<\alpha\leqslant 1$. Clearly this shows that there is always a prime number in the short interval $[x-x^{\frac{1}{2}+\varepsilon},x]$. Unfortunately, we can’t prove (2) unconditionally. Now the best unconditional result is due to Baker, Harman and Pintz. In [1] they showed that $0.525\leqslant\alpha\leqslant 1$ is admissible. But even we assume the Riemann Hypothesis, we can’t extend the range $\frac{1}{2}+\varepsilon<\alpha\leqslant 1$ anymore.
In 2004, under the assumption of the existence of exceptional Dirichlet characters, Friedlander and Iwaniec [3] first extended the range of $\alpha$ to some numbers below $\frac{1}{2}$. Actually they proved the following theorem:
Theorem 1.1.
([[3], Theorem 1.1]). Let $\chi$ be the real primitive character of conductor D. Let $x\geqslant D^{r}$ with $r=18289$ and $\frac{39}{79}\leqslant\alpha\leqslant 1$. Then
$\psi(x)-\psi(x-x^{\alpha})=x^{\alpha}\left\{1+O\left(L(1,\chi)(\log x)^{r^{r}}\right)\right\}$
where $L(s,\chi)$ is the Dirichlet L-function.
In fact, their result strongly depends on a deep result involving product of three Dirichlet $L$-series by themselves ([4], with only classical Van der Corput method):
Theorem 1.2.
([[4], Theorem 4.2]). Let $\chi_{j}\left(\bmod D_{j}\right)$ denote primitive characters and $D=D_{1}D_{2}D_{3}$. For any $x\geqslant 1$, we have
$\Delta:=$ $\sum_{n_{1}n_{2}n_{3}\leqslant x}\chi_{1}\left(n_{1}\right)\chi_{2}\left(n_{2}\right)\chi_{3}\left(n_{3}\right)-\operatorname{Res}_{s=1}\left(L\left(s,\chi_{1}\right)L\left(s,\chi_{2}\right)L\left(s,\chi_{3}\right)\frac{x^{s}}{s}\right)$
$\ll$ $D^{\frac{38}{75}}x^{\frac{37}{75}+\varepsilon}.$
In 2017, based on his previous work on higher order derivative tests for exponential sums [6], Nowak [7] [8] combined the classic Van der Corput method with Huxley’s ”Discrete Hardy–Littlewood method” (see [5]) and obtained the following two estimations of $\Delta$:
Theorem 1.3.
([[7], Theorem 1], [[8], Theorem 3]). Put $D_{\max}:=\max\left(D_{1},D_{2},D_{3}\right)$, then for any $x\geqslant 1$, we have
$\Delta\ll x^{\varepsilon}\left(D^{\frac{1153}{5073}}D_{\max}^{\frac{474}{1691}}x^{\frac{2498}{5073}}+D^{\frac{14077}{81168}}D_{\max}^{\frac{54661}{162336}}x^{\frac{26507}{54112}}+D^{\frac{38017}{144981}}D_{\max}^{\frac{3986}{16109}}x^{\frac{71090}{144981}}+D^{\frac{856679}{2404602}}D_{\max}^{\frac{1281}{6764}}x^{\frac{2185055}{4809204}}\right)$ (A)
and
$\Delta\ll$ $x^{\varepsilon}\left(D^{\frac{1153}{5073}}D_{\max}^{\frac{474}{1691}}x^{\frac{2498}{5073}}+D^{\frac{3043}{12387}}D_{\max}^{\frac{1086}{4129}}x^{\frac{6086}{12387}}+D^{\frac{2719}{13362}}D_{\max}^{\frac{2719}{8908}}x^{\frac{13129}{26724}}+D^{\frac{1}{4}}D_{\max}^{\frac{1}{2}}x^{\frac{1}{4}}\right.$
$\left.+D^{\frac{2}{7}}D_{\max}^{\frac{2}{7}}x^{\frac{3}{7}}+D^{\frac{1}{3}}x^{\frac{1}{3}}+D^{\frac{1}{6}}D_{\max}^{\frac{1}{2}}x^{\frac{1}{3}}+D^{\frac{328}{2719}}D_{\max}^{\frac{1}{2}}x^{\frac{2063}{5438}}+D^{\frac{205}{1422}}D_{\max}^{\frac{1}{2}}x^{\frac{253}{711}}\right).$ (B)
In this paper, we further combine Nowak’s work with Bourgain’s decoupling method (see [2]) and get some similar estimations as Theorem 1.3. We will give a detailed proof of one of our results which is the following one:
Theorem 1.4.
Put $D_{\max}:=\max\left(D_{1},D_{2},D_{3}\right)$, then for any $x\geqslant 1$, we have
$\Delta\ll x^{\varepsilon}\left(D^{\frac{1153}{5073}}D_{\max}^{\frac{474}{1691}}x^{\frac{2498}{5073}}+D^{\frac{24851}{142044}}D_{\max}^{\frac{95873}{284088}}x^{\frac{46171}{94696}}+D^{\frac{38017}{144981}}D_{\max}^{\frac{3986}{16109}}x^{\frac{71090}{144981}}+D^{\frac{856679}{2404602}}D_{\max}^{\frac{1281}{6764}}x^{\frac{2185055}{4809204}}\right).$
Obviously, Theorems 1.31.4 imply the same bound $\Delta\ll D^{\frac{2575}{5073}}x^{\frac{2498}{5073}+\varepsilon}$. Then by the similar arguments as in [3], we can get the following result on primes in short intervals:
Theorem 1.5.
Let $\chi$ be the real primitive character of conductor D. Let $x\geqslant D^{r}$ with $r=1100000$ and $\frac{2500}{5077}\leqslant\alpha\leqslant 1$. Then
$\psi(x)-\psi(x-x^{\alpha})=x^{\alpha}\left\{1+O\left(L(1,\chi)(\log x)^{r^{r}}\right)\right\}.$
remark.
We have $0.49370.4936<\frac{39}{77}<$ 0.4937 and $0.49250.4924<\frac{2500}{5077}<$ 0.4925.
Compare our Theorem 1.5 with Theorem 1.1, we can find that our result is non-trivial if
$L(1,\chi)\ll(\log x)^{-1100000^{1100000}-1}$ (3)
or
$L(1,\chi)\ll(\log D)^{-1100000^{1100000}-1}$ (4)
if $D$ is a positive power of $x$. In fact, Zhang [9] posited that there is no $L$-function with
$L(1,\chi)<c(\log D)^{-2022}$ (5)
for some constant $c$.
2. Higher order derivative tests for single exponential sums
Lemma 2.1.
For $r\geqslant 4$ a fixed integer, and positive real parameters $M\geqslant 1$ and $T$, suppose that $F$ is a real function on some compact interval $I^{*}$ of length $M$, with $r+1$ continuous derivatives satisfying throughout
$F^{(j)}\asymp TM^{-j}\quad\operatorname{for}\quad j=r-2,r-1,r.$
Then, for every interval $I\subseteq I^{*}$,
$E:=\sum_{m\in I}\mathrm{e}^{2\pi\mathrm{i}F(m)}\ll M^{a_{r}}T^{b_{r}}+M^{\xi_{r}}T^{\eta_{r}}+M^{\alpha_{r}}+M^{\gamma_{r}}T^{-\delta_{r}}$
where $a_{4}=\frac{1}{2},b_{4}=\frac{32}{205}+\varepsilon,\xi_{4}=0,\eta_{4}=\frac{31}{84}+\varepsilon,\alpha_{4}=\frac{871}{1086},\gamma_{4}=1,\delta_{4}=\frac{1}{2}$, and for every $j>4$,
$\begin{gathered}a_{j}=\frac{a_{j-1}+b_{j-1}+1}{2\left(b_{j-1}+1\right)},\quad b_{j}=\frac{b_{j-1}}{2\left(b_{j-1}+1\right)},\\ \xi_{j}=\frac{\left(\xi_{j-1}+1\right)\left(b_{j-1}+1\right)-a_{j-1}\eta_{j-1}}{2\left(b_{j-1}+1\right)},\quad\eta_{j}=\frac{\eta_{j-1}}{2\left(b_{j-1}+1\right)},\\ \alpha_{j}=\frac{\alpha_{j-1}+1}{2},\\ \gamma_{j}=\frac{a_{j-1}\delta_{j-1}+\left(b_{j-1}+1\right)\left(\gamma_{j-1}+1\right)}{2\left(b_{j-1}+1\right)},\quad\delta_{j}=\frac{\delta_{j-1}}{2\left(b_{j-1}+1\right)}.\end{gathered}$
Proof.
The proof is very similar to that of [[6], Theorem 1]. The only difference is that we use the bound $E\ll M^{\frac{1}{2}}T^{\frac{13}{84}+\varepsilon}$ in the range $T^{\frac{17}{42}+\varepsilon}\ll M\ll T^{\frac{1}{2}}$. ∎
Lemma 2.2.
For positive real parameters $M\geqslant 1$ and $T$, suppose that $F$ is a real function on some compact interval $I^{*}$ of length $M$, with 6 continuous derivatives satisfying throughout
$F^{(j)}\asymp TM^{-j}\quad\operatorname{for}\quad j=3,4,5.$
Then, for every interval $I\subseteq I^{*}$,
$E\ll M^{\frac{679}{948}}T^{\frac{16}{237}+\varepsilon}+M^{\frac{33461}{79632}}T^{\frac{6355}{39816}+\varepsilon}+M^{\frac{1957}{2172}}+M^{\frac{2101}{1896}}T^{-\frac{205}{948}}.$
Proof.
This is the case $r=5$ of our Lemma 2.1. ∎
3. A problem considered by Friedlander, Iwaniec and Nowak
We follow in all essentials the argument of [4], taking a bit more care and using our Lemma 2.2. Writing $e(w):=\mathrm{e}^{2\pi\mathrm{i}w}$ as usual, we start from the estimate
$\Delta\ll\left|D^{\frac{1}{6}}x^{\frac{1}{3}}\sum_{n_{1}n_{2}n_{3}\leqslant N}\frac{\bar{\chi}_{1}\left(n_{1}\right)\bar{\chi}_{2}\left(n_{2}\right)\bar{\chi}_{3}\left(n_{3}\right)}{\left(n_{1}n_{2}n_{3}\right)^{\frac{2}{3}}}e\left(\pm 3\left(\frac{n_{1}n_{2}n_{3}x}{D}\right)^{\frac{1}{3}}\right)\right|+x^{\varepsilon}\left(\frac{Dx^{2}}{N}\right)^{\frac{1}{3}}$ (6)
where the sign $\pm$ is chosen so that the modulus involved becomes maximal. This result is immediate from [[4], Proposition 3.2]. The exponential sum here can be split up into $O\left(\log^{3}x\right)$ subsums
$S\left(N_{1},N_{2},N_{3}\right):=\sum_{\begin{subarray}{c}n_{j}\sim N_{j}\\ j=1,2,3\end{subarray}}\frac{\bar{\chi}_{1}\left(n_{1}\right)\bar{\chi}_{2}\left(n_{2}\right)\bar{\chi}_{3}\left(n_{3}\right)}{\left(n_{1}n_{2}n_{3}\right)^{\frac{2}{3}}}e\left(\pm 3\left(\frac{n_{1}n_{2}n_{3}x}{D}\right)^{\frac{1}{3}}\right)$
with $N_{1}N_{2}N_{3}\leqslant N$ throughout. Without loss of generality we assume that $N_{1}\leqslant N_{2}\leqslant N_{3}$, and put $P=N_{1}N_{2}N_{3}$ for short. Then obviously
$S\left(N_{1},N_{2},N_{3}\right)\ll P^{-\frac{2}{3}}\sum_{\begin{subarray}{c}n_{j}\sim N_{j}\\ j=1,2\end{subarray}}\left|\sum_{n_{3}\sim N_{3}}\bar{\chi}_{3}\left(n_{3}\right)e\left(\pm 3\left(\frac{n_{1}n_{2}n_{3}x}{D}\right)^{\frac{1}{3}}\right)\right|.$ (7)
Here integration by parts has been used with respect to $n_{3}$, and trivial estimation with respect to $n_{1},n_{2}$. For each fixed pair $\left(n_{1},n_{2}\right)$, the range for $n_{3}$ is chosen in such a way that the absolute value on the right hand side becomes maximal.
The next step is to express the character $\bar{\chi}_{3}$ by means of the Gauss sums $G\left(m,\bar{\chi}_{3}\right)$ :
$\bar{\chi}_{3}\left(n_{3}\right)=\frac{1}{D_{3}}\sum_{m=1}^{D_{3}}G\left(m,\bar{\chi}_{3}\right)e\left(-\frac{mn_{3}}{D_{3}}\right),\quad G\left(m,\bar{\chi}_{3}\right):=\sum_{k=1}^{D_{3}}\bar{\chi}_{3}(k)e\left(\frac{km}{D_{3}}\right)$
Since $\left|G\left(m,\bar{\chi}_{3}\right)\right|\leqslant D_{3}^{\frac{1}{2}}\leqslant D_{\max}^{\frac{1}{2}}$, it follows that for each fixed $\left(n_{1},n_{2}\right)$, with $n_{1}\sim N_{1}$, $n_{2}\sim N_{2}$,
$\sum_{n_{3}\sim N_{3}}\bar{\chi}_{3}\left(n_{3}\right)e\left(\pm 3\left(\frac{n_{1}n_{2}n_{3}x}{D}\right)^{\frac{1}{3}}\right)\ll D_{\max}^{\frac{1}{2}}\max_{m=1,\ldots,D_{3}}\left|\sum_{n_{3}\sim N_{3}}e\left(3\left(\frac{n_{1}n_{2}n_{3}x}{D}\right)^{\frac{1}{3}}\mp\frac{mn_{3}}{D_{3}}\right)\right|.$
The last exponential sums will now be bounded by Lemma 2.2. The conditions of Lemma 2.2 are satisfied with
$M=N_{3},\quad T=\left(\frac{Px}{D}\right)^{\frac{1}{3}}.$
Therefore, by Lemma 2.2 we have
$\sum_{n_{3}\sim N_{3}}\bar{\chi}_{3}\left(n_{3}\right)e\left(3\left(\frac{n_{1}n_{2}n_{3}x}{D}\right)^{\frac{1}{3}}\right)$
$\ll$ $D_{\max}^{\frac{1}{2}}\left(N_{3}^{\frac{679}{948}}\left(\left(\frac{Px}{D}\right)^{\frac{1}{3}}\right)^{\frac{16}{237}+\varepsilon}+N_{3}^{\frac{33461}{79632}}\left(\left(\frac{Px}{D}\right)^{\frac{1}{3}}\right)^{\frac{6355}{39816}+\varepsilon}+N_{3}^{\frac{1957}{2172}}+N_{3}^{\frac{2101}{1896}}\left(\left(\frac{Px}{D}\right)^{\frac{1}{3}}\right)^{-\frac{205}{948}}\right).$ (8)
By (7)–(8) we get that
$D^{\frac{1}{6}}x^{\frac{1}{3}}S\left(N_{1},N_{2},N_{3}\right)$
$\ll$ $D^{\frac{1}{6}}x^{\frac{1}{3}}P^{\frac{1}{3}}D_{\max}^{\frac{1}{2}}\left(\frac{P^{\frac{16}{711}}x^{\frac{16}{711}}}{D^{\frac{16}{711}}N_{3}^{\frac{269}{948}}}+\frac{P^{\frac{6355}{119448}}x^{\frac{6355}{119448}}}{D^{\frac{6355}{119448}}N_{3}^{\frac{46171}{79632}}}+N_{3}^{-\frac{215}{2172}}+\frac{D^{\frac{205}{2844}}N_{3}^{\frac{205}{1896}}}{P^{\frac{205}{2844}}x^{\frac{205}{2844}}}\right)$
$\ll$ $D_{\max}^{\frac{1}{2}}\left(\frac{D^{\frac{205}{1422}}P^{\frac{253}{711}}x^{\frac{253}{711}}}{N_{3}^{\frac{269}{948}}}+\frac{D^{\frac{13553}{119448}}P^{\frac{46171}{119448}}x^{\frac{46171}{119448}}}{N_{3}^{\frac{46171}{79632}}}+\frac{D^{\frac{1}{6}}P^{\frac{1}{3}}x^{\frac{1}{3}}}{N_{3}^{\frac{215}{2172}}}+D^{\frac{679}{2844}}N_{3}^{\frac{205}{1896}}P^{\frac{743}{2844}}x^{\frac{743}{2844}}\right).$ (9)
By the trivial facts $P^{\frac{1}{3}}\leqslant N_{3}\leqslant N$ to get rid of $N_{3}$ and $P\ll N$, we have
$\Delta\ll$ $\left(\frac{Dx^{2}}{N}\right)^{\frac{1}{3}}+D_{\max}^{\frac{1}{2}}\left(D^{\frac{205}{1422}}N^{\frac{743}{2844}}x^{\frac{253}{711}}+D^{\frac{13553}{119448}}N^{\frac{46171}{238896}}x^{\frac{46171}{119448}}+D^{\frac{1}{6}}N^{\frac{1957}{6516}}x^{\frac{1}{3}}+D^{\frac{679}{2844}}N^{\frac{2101}{5688}}x^{\frac{743}{2844}}\right).$ (10)
Finally, by choosing
$N=D^{\frac{538}{1691}}D_{\max}^{-\frac{1422}{1691}}x^{\frac{884}{1691}},$
we obtain the bound
$\Delta\ll$ $x^{\varepsilon}\left(D^{\frac{1153}{5073}}D_{\max}^{\frac{474}{1691}}x^{\frac{2498}{5073}}+D^{\frac{24851}{142044}}D_{\max}^{\frac{95873}{284088}}x^{\frac{46171}{94696}}+D^{\frac{38017}{144981}}D_{\max}^{\frac{3986}{16109}}x^{\frac{71090}{144981}}+D^{\frac{856679}{2404602}}D_{\max}^{\frac{1281}{6764}}x^{\frac{2185055}{4809204}}\right)$
$\ll$ $D^{\frac{2575}{5073}}x^{\frac{2498}{5073}+\varepsilon},$ (11)
which is our Theorem 1.4, implies the same bound as Nowak’s Theorem 1.3(A) but reduces the exponent of $x$ on the second term on the right hand side (we have $0.48990.4898<\frac{26507}{54112}<$ 0.4899 and $0.48760.4875<\frac{46171}{94696}<$ 0.4876).
4. A new application: Primes in Short Intervals
In his articles [6][8], Nowak considered several problems concerning products of $L$-series and pointed that his method is more powerful in the case of exponential sums depending on several parameters. For exponential sums depending on one parameter only, multiple sum estimations are often more powerful. (see [[8], Concluding Remark]). Now we follow Friedlander and Iwaniec directly and show that our Theorem 1.4 can be used to extend the range of $\alpha$. We first introduce some arithmetic functions:
$\lambda(n):=\sum_{ad=n}\chi(d),\quad\nu(m):=\sum_{ad=m}\mu(a)\mu(d)\chi(d),\quad\rho(n):=\sum_{lm=n}\lambda(m),$
and a close relative of $\lambda(d)$, namely
$\lambda^{\prime}(d):=\sum_{kl=d}\chi(k)\log l$
which has the following properties:
$\lambda^{\prime}(d)=\sum_{bc=d}\lambda(b)\Lambda(c),\quad\Lambda(n)=\sum_{dm=n}\lambda^{\prime}(d)\nu(m),\quad 0\leqslant\lambda^{\prime}(d)\leqslant\tau(d)\log d.$
In the classical divisor problem, one need to get an asymptotic formula for the functions:
$\mathcal{D}\left(x;\lambda\right):=\sum_{n\leqslant x}\lambda(n),\quad\mathcal{D}\left(x;\lambda^{\prime}\right):=\sum_{n\leqslant x}\lambda^{\prime}(n)$
and we also need the asymptotic formula for
$\mathcal{D}\left(x;\rho\right):=\sum_{n\leqslant x}\rho(n).$
Lemma 4.1.
For $x\geqslant 1$ we have
$\mathcal{D}\left(x;\lambda\right)=$ $L(1,\chi)x+O\left(D^{\frac{1}{3}}x^{\frac{1}{3}+\varepsilon}\right),$ (12)
$\mathcal{D}\left(x;\lambda^{\prime}\right)=$ $L(1,\chi)x\log x+\left(L^{\prime}(1,\chi)-L(1,\chi)\right)x+O\left(D^{\frac{1}{3}}x^{\frac{1}{3}+\varepsilon}\right),$ (13)
$\mathcal{D}\left(x;\rho\right)=$ $L(1,\chi)x\log x+\left(L^{\prime}(1,\chi)+(2\gamma-1)L(1,\chi)\right)x+O\left(D^{\frac{2575}{5073}}x^{\frac{2498}{5073}+\varepsilon}\right),$ (14)
where $\gamma$ is the Euler constant and the implied constant depends only on $\varepsilon$.
Proof.
These can be proved by similar arguments as in [4], and the only difference is that we use our Theorem 1.4 instead of Theorem 1.2 in order to prove the formula (14). Actually (14) is the Corollary 1.1 in [7]. ∎
Now we split the functions $\rho=\rho^{*}+\rho_{*}$, $\Lambda=\Lambda^{*}+\Lambda_{*}$ and further define
$\rho^{*}(n):=\sum_{\begin{subarray}{c}lm=n\\ m\leqslant D^{2}\end{subarray}}\lambda(m),\quad\rho_{*}(n):=\sum_{\begin{subarray}{c}lm=n\\ m>D^{2}\end{subarray}}\lambda(m),$
$\Lambda^{*}(n):=\sum_{\begin{subarray}{c}dm=n\\ m\leqslant D^{2}\end{subarray}}\lambda^{\prime}(d)\nu(m),\quad\Lambda_{*}(n):=\sum_{\begin{subarray}{c}dm=n\\ m>D^{2}\end{subarray}}\lambda^{\prime}(d)\nu(m),$
$\psi^{*}(x):=\sum_{n\leqslant x}\Lambda^{*}(n),\quad\psi_{*}(x):=\sum_{n\leqslant x}\Lambda_{*}(n).$
Lemma 4.2.
[[3], Corollary 4.2]. For $D^{\frac{5}{2}}x^{\frac{3}{8}}<y\leqslant x$ we have
$\psi^{*}(x)-\psi^{*}(x-y)=y+O\left(L(1,\chi)y(\log x)^{9}\right),$ (15)
where the implied constant is absolute.
By Lemma 4.1 and similar arguments as in [3], for $D^{2}x^{\frac{2499}{5075}}<y\leqslant x$ we obtain the following asymptotic formula
$\mathcal{D}\left(x;\rho_{*}\right)-\mathcal{D}\left(x-y;\rho_{*}\right)=L(1,\chi)y\{\log x+O(1)\}.$ (16)
Now by (16) and
$\left|\psi_{*}(x)-\psi_{*}(x-y)\right|\leqslant(\log x)\sum_{\delta\leqslant\Delta}(2\tau(\delta))^{A}\left[\mathcal{D}\left(x\delta^{-1};\rho_{*}\right)-\mathcal{D}\left((x-y)\delta^{-1};\rho_{*}\right)\right]$ (17)
with the bound
$\sum_{\delta\leqslant\Delta}\tau(\delta)^{A}\delta^{-1}\ll(\log\Delta)^{2^{A}+1}$ (18)
where $\Delta=x^{\frac{1}{r}},A=\frac{r\log r}{\log 2}$ and $r$ is a positive integer, we have
Lemma 4.3.
For $D^{\frac{5}{2}}x^{\frac{2499}{5075}+\frac{2576}{5075r}}<y\leqslant x$ with $r\geqslant 1$, we have
$\psi_{*}(x)-\psi_{*}(x-y)=O\left(L(1,\chi)y(\log x)^{r^{r}+3}\right),$ (19)
where the implied constant depends only on $r$.
Adding (19) to (15) we find that
$\psi(x)-\psi(x-y)=O\left(L(1,\chi)y(\log x)^{r^{r}+3}\right),$ (20)
subject to the conditions of Lemma 4.3. We require $D\leqslant x^{\frac{1}{r}}$ and then choose a number $\theta$ with $\frac{2499}{5075}<\theta<\frac{1}{2}$ satisfies
$\frac{5}{2r}+\frac{2499}{5075}+\frac{2576}{5075r}\leqslant\theta.$
By choosing $\theta=\frac{2500}{5077}$ and $r=1100000$, we complete the proof of Theorem 1.5. We remark that if we only use Bourgain’s result but without using Huxley’s one, we can only obtain the range $0.4931\leqslant\alpha\leqslant 1$. It seems that if we are able to make some modifications, our result may be improved to $0.4923\leqslant\alpha\leqslant 1$, which is rather near to the limit obtained by this method.